OpenStudy (anonymous):

evaluate each expression for the given values of the variables. 16.) k=-1 and n=-2 -ksquared-(3k-5n)+4n

jimthompson5910 (jim_thompson5910):

replace every k with -1 (since k = -1) and replace every n with -2 (since n = -2)

jimthompson5910 (jim_thompson5910):

so -k^2 - (3k - 5n) + 4n -(-1)^2 - (3(-1) - 5(-2)) + 4(-2) evaluate that and simplify

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

the next line after the line u gave was 1^2(-3+10)+-8 next line I got 1(7)+-8 next line I got 7+-8=answer -1 did I msss up anywhere????? those are what I got after your two lines given

jimthompson5910 (jim_thompson5910):

-k^2 - (3k - 5n) + 4n -(-1)^2 - (3(-1) - 5(-2)) + 4(-2) -(1) - (-3 + 10) - 8 -1 - (7) - 8 -1 - 7 - 8 -8 - 8 -16 hopefully you can spot your error

OpenStudy (anonymous):

i didn't catch my error but thanks for the step by step help

OpenStudy (anonymous):

could you help with more please?

jimthompson5910 (jim_thompson5910):

your error is that you had 1(7)+-8 when it should have been -1 - (7) +- 8

OpenStudy (anonymous):

oh ok. thanks wll ou help with more algebra two problems please...????? r u good at story problems i have 6 in a row and don't have the fanest idea how to do them

jimthompson5910 (jim_thompson5910):

and I can help with a few more

jimthompson5910 (jim_thompson5910):

I'll help with a few, but 6 sounds like a lot

jimthompson5910 (jim_thompson5910):

we'll see how it goes

OpenStudy (anonymous):

32.) think about a plan..... Tran's truck gets very poor gas mileage. if tran pays $84 to fill his truck with gas and is able to drive m miles on a fulltank, what expression shows his gas cost per mile?????? what operation does "per" indicate???? check your expression by substituting 200 miles for m. does your answer make sense?????? 33.) suppose you need to subtract a from b but mistakenly subtract b from a instead. how is the answer you get related to the correct answer? explain.

jimthompson5910 (jim_thompson5910):

ok let's focus on #32 first

jimthompson5910 (jim_thompson5910):

it costs $84 to fill up the tank completely and he's able to drive m miles on this full tank so this means that it costs him $84 to drive m miles

jimthompson5910 (jim_thompson5910):

the word "per" means "divide" eg: miles per hour = miles/hour

jimthompson5910 (jim_thompson5910):

so when they say "cost per mile" they mean cost/mile = $84/m = 84/m

OpenStudy (anonymous):

i only have 4 story problems not 6

jimthompson5910 (jim_thompson5910):

so the cost per mile is 84/m

jimthompson5910 (jim_thompson5910):

so for instance, if he can drive 84 miles on a full tank, then m = 84 and 84/m = 84/84 = 1 which means that the cost per mile is $1 per mile, ie it costs $1 to drive 1 mile

jimthompson5910 (jim_thompson5910):

what happens when m = 200?

OpenStudy (anonymous):

84/200?

jimthompson5910 (jim_thompson5910):

which reduces to what

OpenStudy (anonymous):

50/21?

jimthompson5910 (jim_thompson5910):

you have it flipped and it should be 21/50

jimthompson5910 (jim_thompson5910):

now convert that to decimal form

OpenStudy (anonymous):

your going really fast. and all the messages at once are confusing. idk what to write for my work or answers on 32 sorry but im a slow learner. im a lil lost.. can u start over explaining one message at a time so i get it???????????????/

jimthompson5910 (jim_thompson5910):

you saw how I got 84/m right?

OpenStudy (anonymous):

84 to fill up tank of gas. m stands for number of miles driven. do i write down 84/m irst on my paper?

jimthompson5910 (jim_thompson5910):

that's the expression for the cost per mile, so yes

OpenStudy (anonymous):

k got it

jimthompson5910 (jim_thompson5910):

now you replace m with 200 to get 84/m = 84/200 = 21/50 = 0.42 you use a calculator to convert to decimal form (or you can use long division) this means that it costs him $0.42 or 42 cents to drive one mile so if he can go 200 miles on a full tank of gas, then it costs 42 cents per mile

OpenStudy (anonymous):

whered you get 84/200? 21/50? 0.42?

jimthompson5910 (jim_thompson5910):

in the instructions, it says to use m = 200 it says "check your expression by substituting 200 miles for m"

OpenStudy (anonymous):

ok got it

jimthompson5910 (jim_thompson5910):

84/200 then reduces to 21/50 you use a calculator to get 21/50 = 0.42

OpenStudy (anonymous):

21/50 you reduced to get thatoooookk think i ot it

OpenStudy (anonymous):

thanks!! now, what would i exactly circle for my answers? it had several answer questions

jimthompson5910 (jim_thompson5910):

you would circle 84/m and 0.42 oh and don't forget to say whether the answer of 0.42 makes sense or not

OpenStudy (anonymous):

i have four circles. 84/m. 0.42. makes sense. per means divide.. am i right only four circles?

jimthompson5910 (jim_thompson5910):

so far I'd say 2 or 3 for #32

OpenStudy (anonymous):

oh wait! what expression shows his gas cost per mile???

jimthompson5910 (jim_thompson5910):

84/m

OpenStudy (anonymous):

do i have every answer circled? i said two messages up what i circled

jimthompson5910 (jim_thompson5910):

yes it sounds like you got everything covered

OpenStudy (anonymous):

thank u soooo much for your help sir.33? 34? 35? got time to help with 3 or a couple after that more? I REALLY do appreciate it (=

jimthompson5910 (jim_thompson5910):

sure we can do 3 more

OpenStudy (anonymous):

33 i typed above with 32 already

jimthompson5910 (jim_thompson5910):

33.) suppose you need to subtract a from b but mistakenly subtract b from a instead. how is the answer you get related to the correct answer? explain.

OpenStudy (anonymous):

THANK YOU SOOOO MUCH sir(; your my savior

jimthompson5910 (jim_thompson5910):

it's best to do some examples

jimthompson5910 (jim_thompson5910):

pick any two different numbers you want

OpenStudy (anonymous):

examples then you will help with my homework?????

jimthompson5910 (jim_thompson5910):

this is helping, it will help you see how to answer

jimthompson5910 (jim_thompson5910):

go ahead and pick any two numbers you want

OpenStudy (anonymous):

34 and 35 ok? they are story problems?

jimthompson5910 (jim_thompson5910):

no not problem numbers, I meant any two numbers you can think of for #33

OpenStudy (anonymous):

2 and -4

jimthompson5910 (jim_thompson5910):

ok for now, let's keep them positive, so let's do 2 and 4

jimthompson5910 (jim_thompson5910):

let a = 2 and b = 4

jimthompson5910 (jim_thompson5910):

if a = 2 and b = 4 then what is a - b equal to?

OpenStudy (anonymous):

-2?

jimthompson5910 (jim_thompson5910):

good, now what is b - a using the same values for 'a' and 'b'

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

so to recap a-b = -2 b-a = 2

OpenStudy (anonymous):

ya?

jimthompson5910 (jim_thompson5910):

let's pick another pair of numbers, say 7 and 10 if a = 7 and b = 10, then a-b = -3 b-a = 3

jimthompson5910 (jim_thompson5910):

see what's going on?

OpenStudy (anonymous):

yes i do your picking easy ones

jimthompson5910 (jim_thompson5910):

so in general, what's the connection between a-b and b-a

OpenStudy (anonymous):

its the same answer. just when you take b-a it equals a negative integer

OpenStudy (anonymous):

instead of a positive integer

jimthompson5910 (jim_thompson5910):

good, so if a-b is a negative number then b-a is the positive version of that number OR if a-b is a positive number, then b-a is the negative version of that same number

OpenStudy (anonymous):

that's the answer yay! we are two smart cookies teamwork

OpenStudy (anonymous):

right? your last message is the answer?

jimthompson5910 (jim_thompson5910):

lol so if you subtract in the wrong order and you know the answer has to be positive, then you can just drop the negative sign

jimthompson5910 (jim_thompson5910):

yes pretty much

jimthompson5910 (jim_thompson5910):

another way to put the answer: the two expressions a-b and b-a are equal in absolute value since they both produce the same result if you ignore the positive and negative signs

OpenStudy (anonymous):

that answer sounds super smart your last message thnks a billion. but PLEASE don't leave yet

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

for numbers 34 and 35 write an algebraic expression to model each situation. Thank u soooo much for your help btw i really do appreciate it

OpenStudy (anonymous):

34.)jobs... you have a summer job at a car wash. you earn $8.50 per hour and are expected to pay a one-time fe of $15 for he unifom. If you work x hours per week, how much will you make durng the first week????????????? 35.)class project.. the frshman class will be selling carnations as a class project. what is their income after they pay the florist a flat fee of $200 and sell x carnations for $2 each. name the property of real numbers illustrated by the equation 44.)2(s-t)=2s-2t 46.) -(2t-11)=11-2t simplify each expression 56.) 4x+3x-4 58.) -2(4+b)+4(b-5)

jimthompson5910 (jim_thompson5910):

were you able to get anywhere with #34

OpenStudy (anonymous):

$8.50ph+$15=xh per week=? something like that?

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

if you work for x hours per week and you earn $8.50 an hour, then you earn 8.50x dollars for working those x hours subtract off the 15 dollars you have to pay for the uniform to get 8.50x - 15

jimthompson5910 (jim_thompson5910):

so you will make 8.50x - 15 dollars

jimthompson5910 (jim_thompson5910):

if it told you how many hours a week you worked that first week, then you could find an actual number for the answer

OpenStudy (anonymous):

there isn't a actual answer number? just the algebraic expression?

jimthompson5910 (jim_thompson5910):

yes because it doesn't say how many hours as week you worked

jimthompson5910 (jim_thompson5910):

it just says "x hours"

OpenStudy (anonymous):

i get it how u explained it. i wish u was my teacher whoever u may be

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

if it said you worked 40 hours that first week, then you'd replace x with 40 and evaluate

OpenStudy (anonymous):

how many $ signs are in the answer 2 isn't it?

jimthompson5910 (jim_thompson5910):

what do you mean

jimthompson5910 (jim_thompson5910):

I think you mean decimal places right?

OpenStudy (anonymous):

no $ signs dollar signs